# Surrogate Modeling of Landau Damping with Deep Operator Networks

### v1

@@ 1 Introduction
In space and astrophysical plasmas, most physical processes are collisionless. Generally, simulating these collisionless processes demands kinetic approaches, such as the particle-in-cell (PIC) or Vlasov methods. However, these kinetic simulations are computationally demanding and therefore not well-suited for efficiently addressing large-scale problems that involve collisionless physics. To manage this challenge with affordable computational costs, two main approaches have been developed for large-scale global simulations: the magnetohydrodynamics with embedded particle-in-cell (MHD-EPIC) model (L. K. S. Daldorff et al., 2014; G. Tóth et al., 2016; Y. Chen et al., 2017; H. Zhou et al., 2020) and the multi-moment multi-fluid model (L. Wang et al., 2015, 2018; C. Dong et al., 2019; L. Wang et al., 2020; S. Jarmak et al., 2020; E. Rulke et al., 2021). Recently, the MHD-EPIC model has evolved to include adaptively embedded PIC regions (MHD-AEPIC) (X. Wang et al., 2022; Y. Chen et al., 2023; D. Li et al., 2023), enabling flexibility to capture localized regions where kinetic effects are significant. On the other hand, substantial efforts have been dedicated to integrating kinetic effects into the multi-moment multi-fluid model through machine learning techniques (C. Ma et al., 2020; E. P. Alves & F. Fiuza, 2022; W. Cheng et al., 2023; Y. Qin et al., 2023; J. Donaghy & K. Germaschewski, 2023; Z. Huang et al., 2025; E. R. Ingelsten et al., 2025).
@@
DeepONets: The foundation of DeepONets is based on the universal approximation theorem (K. Hornik et al., 1989), which guarantees that neural networks can approximate any continuous function with arbitrary accuracy. DeepONets designed to learn mappings between function spaces, making them useful in the context of partial differential equations (PDEs).
As a reference, consider the parametric PDEs of the form:
@@
where $\mathcal{N}$ can be a linear or nonlinear differential operator, $T$ denotes the parametric input and $\mathcal{E}$ is the corresponding functional output. DeepONets can be used to capture the relationship between $T$ and $\mathcal{E}$, represented as
@@
where $t$ denotes the collocation points, $\omega$ represents the network parameters, and $b_{k}$ and $\tau_{k}$ are the outputs of branch and trunk networks as shown in Figure 1. The function $T$ evaluated at fixed sensors $\{\eta_{i}\}_{i=1}^{m}$ will be used as the input of the branch network. For the Landau damping problem, $G_{\omega}$ maps a set of temperature values, $T$, to the electric field energy, $G_{\omega}(T)(t)$, at different time steps, $t$.
Figure 1: The DeepONets architecture consists of two fully-connected neural networks (FNN): the branch network, which encodes temperature values, $T$, and the trunk network, which encodes the coordinates (here the coordinates are time, $t$). The outputs of these networks are combined to approximate the electric field energy. The final output is $G_{\omega}(T)(t)$, where $\omega$ represents the model’s learned weights and $T$ is a vector as $\{T_{i}\}_{i=1}^{N}$.
For different kinds of problems, the branch and trunk networks can be residual network (ResNet), convolutional neural network (CNN), recurrent neural network (RNN), or feed forward neural network (FFNN). In high-dimensional problems where $t$ is a vector with $d$ components, the dimension of $t$ no longer matches the dimension of $T(\eta_{i})$ for $i=1,2,\ldots,m$, and at least two sub-networks are necessary to handle $[T(\eta_{1}),T(\eta_{2}),\ldots,T(\eta_{m})]^{T}$ and $t$ separately. Furthermore, depending on the number and characteristics of input functions, one can incorporate multiple branch networks instead of one (L. Lu et al., 2021).
@@ 3.1 Physical model for data generation
where $f_{s}(x,\textbf{v}_{\textbf{s}},t)$ is the velocity distribution function, $\frac{e_{s}}{m_{s}}$ is the charge-to-mass ratio of the species $s$. Meanwhile, the Poisson equation governs the electrostatic potential generated by charge distributions in a plasma.
@@
where $\phi(x,t)$ is the electrostatic potential, and $\varepsilon_{0}$ is the vacuum permitivity. $\rho(x,t)=\sum_{s}e_{s}n_{s}$ is the charge density, where $e_{s}$ is the charge and
[…]
is number density of the particle species $s$.
@@
In the present work, we use DeepONets to predict the dynamical evolution of electric field energy and thus the rate of Landau damping.
[…]
@@ 3.2 Data for training and testing
where $n_{0}$ is the initial uniform number density of ions and electrons without perturbation. $A_{i}$ and $k_{i}$ are the amplitude and wave number of each perturbed mode, respectively. The dispersion relation of the least damped mode in a uniform, electrostatic plasma with immobile ion background and temperature T =1 is shown in Figure 2, where the left and right panels depict the oscillatory (real) frequency, $\omega_{R}$, and damping rate ($\gamma$; negative growth rate), respectively, versus the wavenumber $k$. Normalized values are used for simplicity. As can be clearly seen, in the long wavelength limit ($k\ll\lambda_{e}^{-1}$), there is little damping. As the wavenumber grows past about $k\sim 0.2\lambda_{e}^{-1}$, the damping rate increases rapidly, indicating that short-wavelength modes are damped more quickly. For our study, in the single-mode case, we use $k=0.35$ (and a small perturbation $A=0.05$) as marked by the blue dot in the right panel of Figure 2. This serves as a convenient baseline case where the damping is substantial but not too aggressive. For the five-mode case, several wavenumbers $k_{i}$ are chosen around this baseline value to include modes with relatively comparable damping rates to allow meaningful competition between each other. The $k_{i}$ values and the corresponding perturbation magnitudes, $A_{i}$, are described in Table. 1.
@@
Table 1: Wavenumber and perturbation amplitude of each mode.
@@ 4 Results and Discussion
In this section, we demonstrate the capability of DeepONets for two different cases. By using DeepONets, the inputs of the branch net are temperature values which are randomly chosen in the range $[0.5,1.5]$, and the inputs of the trunk net are equidistant points $t^{j}=j\Delta t\ (j=0,1,\cdots,J)$ with a step size of $\Delta t$. The output of the network is the electric field energy. During the training phase, the model is optimized by minimizing the loss function to determine the weights. In the testing phase, the trained model is applied to unseen data to assess its accuracy and generalization capability. The network architecture and hyperparameters have been specified in Table 2, and the optimization algorithm employed is the Adaptive Moment Estimation (Adam) method, with an exponential learning rate decay starting at $0.001$.
Table 2: DeepONets architectures and hyperparameters used for training and testing.
@@
Single mode $200$ $50$ $6$ $200$ $\tanh$ Adam $10^{6}$
Five modes $400$ $100$ $6$ $200$ $\tanh$ Adam $10^{6}$
@@ 4.1 Single-mode case
[…]
@@
Table 3: Statistics of relative $L^{2}$ errors for the single-mode case.
@@
The loss function used is the mean square error (MSE), which is defined as follows,
$\mathrm{MSE}=\frac{1}{N}\sum_{i=1}^{N}(\mathcal{E}(t)-G_{\omega}(T_{i})(t))^{2}.$ (10)
$N$ is the number of samples, $G_{\omega}(T_{i})(t)$ is the predicted value corresponding to $T_{i}$, and $\mathcal{E}(t)$ is the true value. In the loss function, $\omega$ is used because it represents the model’s trainable parameters that are adjusted to minimize the error between the predicted and the true values. This allows the model to learn and improve its predictions during the training process.
4.2 Five-modes case
Now, let us consider the initial condition consists of five modes, with the corresponding values listed in Table 1. The network is trained using $400$ samples over the interval $[0,40\omega_{pe}^{-1}]$ with a time step of $\Delta t=0.002\omega_{pe}^{-1}$, and the test set consists of $100$ samples.
[…]
Figure 5 compares the DeepONets predictions with the corresponding reference solutions at different temperature values ($T$). As in the single-mode case, the test cases shown were not included during the training phase, highlighting the model’s generalization capability. Given the relatively large wave perturbation amplitudes $A_{n}$ listed in Table 1 for certain modes, nonlinear Landau damping is expected to occur. In the nonlinear regime (e.g., see the top-left panel of Figure 5), large perturbations drive significant energy transfer from the wave to resonant particles moving slightly below the wave phase speed, accelerating them beyond it. These faster particles subsequently transfer energy back to the wave, leading to an oscillatory exchange of energy. Nonlinear Landau damping is characteristically accompanied by the formation of electron phase-space holes (see, e.g., Figure 4 in Z. Huang et al., 2025). The comparison again demonstrates that the DeepONets surrogate model accurately captures the complex plasma dynamics, with its predictions in close agreement with the reference solutions.
Table 4: Mean relative $L^{2}$ error norms for the five-mode case with different numbers of training and test samples.
@@ 4.2 Five-modes case
The convergence of the loss functions with different numbers of training samples, depicted in Figure 6, tends towards smaller values of $10^{-4}$. This demonstrates that the DeepONets surrogate model not only effectively learns from the training data but also improves its accuracy on test data as the size of the training dataset increases, which again highlights the model’s robustness and generalization capability.
@@ 5 Conclusion
This work was partially supported by NASA grant 80NSSC23K0908, DOE grant DE-SC0024639, and the Alfred P. Sloan Research Fellowship. The authors thank Liang Wang for insightful discussions and for providing the dispersion relation calculations. We would like to acknowledge high-performance computing support from the NASA High-End Computing Program through the NASA Advanced Super-computing Division at Ames Research Center, from National Energy Research Scientific Computing Center, a DOE Office of Science user facility, and from the Derecho system (doi:10.5065/qx9a-pg09) provided by the NSF National Center for Atmospheric Research (NCAR), sponsored by the National Science Foundation. For distribution of the model results used in this study, please contact the corresponding authors. {contribution} C.D. supervised the project and, together with S.S., designed the research. S.S. developed the machine learning architecture based on the DeepONet model and completed the data analyses. Z.H. generated the simulation data. S.S. and C.D. wrote the manuscript. All authors reviewed and approved the final version.
@@ Appendix A Error Analysis and Convergence Behavior in the Five-Mode Case
Figure 7: Five-mode case: The left panel shows the error norms for training and testing as a function of the number of training samples. The right panel depicts the training and test losses versus the number of training samples. The error norms and losses are calculated using Equations (9) and (10).
Figure 8: Five-mode case: Histogram (left) and scatter plot (right) of the relative $L^{2}$ errors for the case with 800 training samples and 200 test samples.

### v2

@@ 1 Introduction
In space and astrophysical plasmas, most physical processes are collisionless. Generally, simulating these collisionless processes demands kinetic approaches, such as the particle-in-cell (PIC) or Vlasov methods. However, these kinetic simulations are computationally demanding and therefore not well suited for efficiently addressing large-scale problems that involve collisionless physics. To manage this challenge with affordable computational costs, two main approaches have been developed for large-scale global simulations: the magnetohydrodynamics with embedded particle-in-cell (MHD-EPIC) model (L. K. S. Daldorff et al., 2014; G. Tóth et al., 2016; Y. Chen et al., 2017; H. Zhou et al., 2020) and the multi-moment multi-fluid model (L. Wang et al., 2015, 2018; C. Dong et al., 2019; L. Wang et al., 2020; S. Jarmak et al., 2020; E. Rulke et al., 2021). Recently, the MHD-EPIC model has evolved to include adaptively embedded PIC regions (MHD-AEPIC) (X. Wang et al., 2022; Y. Chen et al., 2023; D. Li et al., 2023), enabling flexibility to capture localized regions where kinetic effects are significant. On the other hand, substantial efforts have been dedicated to integrating kinetic effects into the multi-moment multi-fluid model through machine learning techniques (C. Ma et al., 2020; E. P. Alves & F. Fiuza, 2022; W. Cheng et al., 2023; Y. Qin et al., 2023; J. Donaghy & K. Germaschewski, 2023; Z. Huang et al., 2025; E. R. Ingelsten et al., 2025).
@@
DeepONets: The foundation of DeepONets is based on the universal approximation theorem (K. Hornik et al., 1989), which guarantees that neural networks can approximate any continuous function with arbitrary accuracy. DeepONets are designed to learn mappings between function spaces, making them useful in the context of partial differential equations (PDEs).
As a reference, consider the parametric PDEs of the form
@@
where $\mathcal{N}$ can be a linear or nonlinear differential operator, $T$ denotes the parametric input, and $\mathcal{E}$ is the corresponding functional output. DeepONets can be used to capture the relationship between $T$ and $\mathcal{E}$, represented as
@@
where $t$ denotes the collocation points, $\omega$ represents the network parameters, and $b_{k}$ and $\tau_{k}$ are the outputs of branch and trunk networks, as shown in Figure 1. The function $T$ evaluated at fixed sensors $\{\eta_{i}\}_{i=1}^{m}$ will be used as the input of the branch network. For the Landau damping problem, $G_{\omega}$ maps a set of temperature values, $T$, to the electric field energy, $G_{\omega}(T)(t)$, at different time steps, $t$.
Figure 1: The DeepONets architecture consists of two fully connected neural networks (FNN): the branch network, which encodes temperature values, $T$, and the trunk network, which encodes the coordinates (here the coordinates are time, $t$). The outputs of these networks are combined to approximate the electric field energy, $\int|E_{x}|^{2}dx$. The final output is $G_{\omega}(T)(t)$, where $\omega$ represents the model’s learned weights and $T$ is a vector as $\{T_{i}\}_{i=1}^{N}$.
For different kinds of problems, the branch and trunk networks can be residual network (ResNet), convolutional neural network (CNN), recurrent neural network (RNN), or feed forward neural network (FFNN). In high-dimensional problems, where $t$ is a vector with $d$ components, the dimension of $t$ no longer matches the dimension of $T(\eta_{i})$ for $i=1,2,\ldots,m$, and at least two sub-networks are necessary to handle $T(\eta_{1}),T(\eta_{2}),\ldots,T(\eta_{m})]^{T}$ and $t$ separately. Furthermore, depending on the number and characteristics of input functions, one can incorporate multiple branch networks instead of one (L. Lu et al., 2021).
@@ 3.1 Physical Model for Data Generation
where $f_{s}(x,v_{s},t)$ is the velocity distribution function and $\frac{e_{s}}{m_{s}}$ is the charge-to-mass ratio of the species, $s$. Meanwhile, the Poisson equation governs the electrostatic potential generated by charge distributions in a plasma:
@@
where $\phi(x,t)$ is the electrostatic potential and $\varepsilon_{0}$ is the vacuum permitivity. $\rho(x,t)=\sum_{s}e_{s}n_{s}$ is the charge density, where $e_{s}$ is the charge and
$n_{s}(x,t)=\int f_{s}(x,v_{s},t)\ dv_{s},$ (5)
is number density of the particle species, $s$.
@@
In the present work, we use DeepONets to predict the dynamical evolution of electric field energy and thus the rate of Landau damping:
$\mathcal{E}(t)=\int|E_{x}(x,t)|^{2}dx$ (6)
@@ 3.2 Data for Training and Testing
where $n_{0}$ is the initial uniform number density of ions and electrons without perturbation. $A_{i}$ and $k_{i}$ are the amplitude and wavenumber of each perturbed mode, respectively. The dispersion relation of the least damped mode in a uniform, electrostatic plasma with immobile ion background and temperature $T=1$ is shown in Figure 2, where the left and right panels depict the oscillatory (real) frequency, $\omega_{R}$, and damping rate ($\gamma$; negative growth rate), respectively, versus the wavenumber, $k$. Normalized values are used for simplicity. As can be clearly seen, in the long wavelength limit ($k\ll\lambda_{e}^{-1}$), there is little damping. As the wavenumber grows past about $k\sim 0.2\lambda_{e}^{-1}$, the damping rate increases rapidly, indicating that short-wavelength modes are damped more quickly. For our study, in the single-mode case, we use $k=0.35$ (and a small perturbation $A=0.05$) as marked by the blue dot in the right panel of Figure 2. This serves as a convenient baseline case where the damping is substantial but not too aggressive. For the five-mode case, several wavenumbers $k_{i}$ are chosen around this baseline value to include modes with relatively comparable damping rates to allow meaningful competition between each other. The $k_{i}$ values and the corresponding perturbation magnitudes, $A_{i}$, are described in Table. 1.
@@
Table 1: Wavenumber and Perturbation Amplitude of Each Mode
@@ 4 Results and Discussion
In this section, we demonstrate the capability of DeepONets for two different cases. By using DeepONets, the inputs of the branch net are temperature values that are randomly chosen in the range $[0.5,1.5]$, and the inputs of the trunk net are equidistant points $t^{j}=j\Delta t\ (j=0,1,\cdots,J)$ with a step size of $\Delta t$. The output of the network is the electric field energy. During the training phase, the model is optimized by minimizing the loss function to determine the weights. In the testing phase, the trained model is applied to unseen data to assess its accuracy and generalization capability. The network architecture and hyperparameters have been specified in Table 2, and the optimization algorithm employed is the Adaptive Moment Estimation (Adam) method, with an exponential learning rate decay starting at $0.001$.
Table 2: DeepONets Architectures and Hyperparameters Used for Training and Testing
@@
Single mode $200$ $50$ 66 $200$ $\tanh$ Adam $10^{6}$
Five modes $400$ $100$ 66 $200$ $\tanh$ Adam $10^{6}$
@@ 4.1 Single-mode Case
Figure 3: Single-mode case: comparison of DeepONets predictions with reference solutions obtained from the Gkeyll Vlasov model for a range of temperature values ($T$). The vertical axis shows the electric field energy, $\int|E_{x}|^{2}dx$, on a logarithmic scale, and the horizontal axis represents time in $\omega_{pe}^{-1}$.
@@
Table 3: Statistics of Relative $L^{2}$ Errors for the Single-mode Case
@@
The loss function used is the mean square error (MSE), which is defined as follows:
$\mathrm{MSE}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{J}(\mathcal{E}_{i}(t^{j})-G_{\omega}(T_{i})(t^{j}))^{2}.$ (10)
NN is the number of samples, $G_{\omega}(T_{i})(t^{j})$ is the predicted value corresponding to $T_{i}$ at $t^{j}$, and $\mathcal{E}_{i}(t^{j})$ is the true value. In the loss function, $\omega$ is used because it represents the model’s trainable parameters that are adjusted to minimize the error between the predicted and the true values. This allows the model to learn and improve its predictions during the training process.
4.2 Five-mode Case
Now, let us consider the initial condition consists of five modes, with the corresponding values listed in Table 1. The network is trained using $400$ samples over the interval $0,40\omega_{pe}^{-1}]$ with a time step of $\Delta t=0.002\omega_{pe}^{-1}$, and the test set consists of $100$ samples.
Figure 5: Five-mode case: comparison between DeepONet predictions and reference Gkeyll Vlasov solutions for varying temperatures ($T$). The vertical axis shows the electric field energy, $\int|E_{x}|^{2}dx$, on a logarithmic scale, and the horizontal axis represents time in $\omega_{pe}^{-1}$.
Figure 5 compares the DeepONets predictions with the corresponding reference solutions at different temperature values ($T$). As in the single-mode case, the test cases shown were not included during the training phase, highlighting the model’s generalization capability. Given the relatively large wave perturbation amplitudes, $A_{n}$ listed in Table 1 for certain modes, nonlinear Landau damping is expected to occur. In the nonlinear regime (e.g., see the top-left panel of Figure 5), large perturbations drive significant energy transfer from the wave to resonant particles moving slightly below the wave phase speed, accelerating them beyond it. These faster particles subsequently transfer energy back to the wave, leading to an oscillatory exchange of energy. Nonlinear Landau damping is characteristically accompanied by the formation of electron phase-space holes (see, e.g., Figure 4 in Z. Huang et al., 2025). The comparison again demonstrates that the DeepONets surrogate model accurately captures the complex plasma dynamics, with its predictions in close agreement with the reference solutions.
Table 4: Mean Relative $L^{2}$ Error Norms for the Five-mode Case with Different Numbers of Training and Test Samples
@@ 4.2 Five-mode Case
The convergence of the loss functions with different numbers of training samples, depicted in Figure 6, tends toward smaller values of $10^{-4}$. This demonstrates that the DeepONets surrogate model not only effectively learns from the training data but also improves its accuracy on test data as the size of the training dataset increases, which again highlights the model’s robustness and generalization capability.
@@ 5 Conclusion
This work was partially supported by NASA grant 80NSSC23K0908, DOE grant DE-SC0024639, and the Alfred P. Sloan Research Fellowship. The authors thank Liang Wang for insightful discussions and for providing the dispersion relation calculations. We would like to acknowledge high-performance computing support from the NASA High-End Computing Program through the NASA Advanced Super-computing Division at Ames Research Center, from National Energy Research Scientific Computing Center, a DOE Office of Science user facility, and from the Derecho system (doi:10.5065/qx9a-pg09) provided by the NSF National Center for Atmospheric Research (NCAR), sponsored by the National Science Foundation. For distribution of the model results used in this study, please contact the corresponding authors. {contribution} C.D. supervised the project and, together with $S.$, designed the research. $S.$ developed the machine learning architecture based on the DeepONet model and completed the data analyses. Z.H. generated the simulation data. $S.$ and C.D. wrote the manuscript. All authors reviewed and approved the final version.
@@ Appendix A Error Analysis and Convergence Behavior in the Five-Mode Case
Figure 7: Five-mode case: the left panel shows the error norms for training and testing as a function of the number of training samples. The right panel depicts the training and test losses vs. the number of training samples. The error norms and losses are calculated using Equations (9) and (10).
Figure 8: Five-mode case: histogram (left) and scatter plot (right) of the relative $L^{2}$ test errors for the case with 800 training samples and 200 test samples.